[Find the point that was used in this point-slope form equation:] y+4=7(x-7)
Answer by TimothyLamb(4379)

Get the equation in point-slope form: y - y1 = m(x - x1)
---
y + 4 = 7(x - 7)
y - (-4) = 7(x - 7)
---
The point used is:
p1 = (7, -4)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php